Three forces acting on an object are given F_1 = (
Answer #1

1.

Fnet = F1+F2+F3 = (-2.45 i -2.40j) + (5.40 i) + (43.5 i + 47.0 j) = (46.45i + 50.0j) N

a) θ=tan^-1(Fnety/Fnetx)= tan^-1(50.0/46.45) = 47.11 deg …above x axis

b) Fnet = sqrt(Fnetx^2+Fnety^2)= sqrt(46.45^2+50.0^2) = 68.25 N

m= Fnet/anet = 68.25/3.50 = 19.5 kg

c) vf= vi+at = 0 + 3.50*18.0 = 63.0 m/s

d) vfx = vix + anetx*t = 0 + 3.50cos47.11*18.0 = 42.88 m/s i

vfy = viy + anety*t = 0 + 3.50sin47.11*18.0 = 46.16 m/s j

vf = (42.88 m/s) i + (46.16 m/s) j

2.

Applying Newton’s second law at point B,

T3=mg = 500 N

Now applying Newton’s second law at point A horizontally,

T2cos27 - T1cos50 =0

T2= T1*(cos50/cos27) -------(1)

Now applying Newton’s second law at point A vertically,

T2sin27 + T1sin50 = T3----------(2)

From (1),

[T1*(cos50/cos27)]sin27 + T1sin50 = T3

T1*(cos50*tan27) + T1sin50 = T3

T1[(cos50*tan27) + sin50] = T3

T1[(cos50*tan27) + sin50] = 500

T1= 457.22 N